OpenStudy (anonymous):

\frac{ 4 }{ x+5 }+\frac{ 5 }{ x^2-5x+25 }-\frac{ 300 }{ x^3+125 } ?

OpenStudy (anonymous):

\[\frac{ 4 }{ x+5 }+\frac{ 5 }{ x^2-5x+25 }-\frac{ 300 }{ x^3+125 }\]

jagr2713 (jagr2713):

is there an answer choice

jagr2713 (jagr2713):

we simplify or factor?

OpenStudy (anonymous):

simplify

OpenStudy (imstuck):

you simplfy by factoring. Do it one fraction at a time. Or else you'll go nuts.

OpenStudy (imstuck):

The first one is as simple as it's going to get, so lets move on the next one, which looks pretty hairy.

OpenStudy (imstuck):

The middle one's denomintor doesn't factor. You get an imaginary number. You're going to be in trouble with this one. It's hugely long.

OpenStudy (anonymous):

thanks lol